The hypotenuse of a right triangle is 5 inches long. The longer leg is 1 inch longer than the shorter leg. Find the side lengths of the triangle.
Answer by ikleyn(52788)

.
The sides are x, (x+1) and 5 (hypotenuse).


x^2 + (x+1)^2 = 5^2


2x^2 + 2x + 1 = 25


2x^2 + 2x - 24 = 0


x^2 + x - 12 = 0


(x+4)*(x-3) = 0


The only positive solution is x= 3.


ANswer.  The legs are 3 and 4 inches.  Classic (3,4,5)-right angled triangle.